OpenStudy (anonymous):

Prime or composite? 27 & 29

OpenStudy (anonymous):

do you know what a prime number is

OpenStudy (anonymous):

prime is 2,3,5,7 and so on

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Heyloo??

OpenStudy (unklerhaukus):

you can write 27 as a product of two primes

OpenStudy (unklerhaukus):

*three

OpenStudy (anonymous):

no i mean is 29&27 prime or composite

OpenStudy (anonymous):

nvm i got it 29 prime 27 composite :)

OpenStudy (anonymous):

A prime number is a natural number that can be divided, without leaving any remainder, by only itself and one. A prime number has only two factors, itself and one. For example, 5 can be divided, without a remainder, only by 5 and 1. 5 has exactly two natural number factors, 5 and 1. 5 is a prime number.

OpenStudy (anonymous):

29 prime, 27 composite

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

A composite number is an natural number that can be divided, without leaving any remainder, by a natural number other than itself and one. For example, 2 = 1 × 2 Prime 3 = 1 × 3 Prime 6 = 1 × 6 and 2 × 3 Composite 6 can be divided by 2 and by 3, so 6 is composite. 15 = 1 × 15 and 3 × 5 Composite 15 can be divided by 3 and by 5, so 15 is composite.